Solve (4x + 3)^2 = 5
Answer by dabanfield(803)

Solve (4x + 3)^2 = 5
Rewrite as:
(4x+3) * (4x+3) = 5
Expand:
16*(x^2) + 24*x + 9 = 5
16*x^2 + 24*x + 4 = 0
(2x+1)*(2x+1) = 0
Solve for x.